Nicole:

If a line segment is 10 inches long, the segment bisector makes two 5 inch long segments. True False

Nicole:

@Gdeinward

Gdeinward:

ah we meet again

Gdeinward:

what do you think?

Nicole:

I think true

Gdeinward:

That is correct. GOOD JOB

Nicole:

Okay ty A segment bisector intersects a segment at its midpoint. True False true?

Gdeinward:

TRUE. GOOD JOB

Nicole:

The distance on the number line from 5 to 12 equals 3x – 2. What is the value of x?

Gdeinward:

Can you post it a little differently? I dont understand

Nicole:

that is the question. there is no graphs or anything

Nicole:

1 attachment
Gdeinward:

I guess you should subtract 5 from 12 and solve for x. So 7=3x-2

Gdeinward:

so 9=3x x = 3? I dunno

Nicole:

So you're not positive about it?

Nicole:

I think thats right what do you think @mhchen

mhchen:

This question is extremely hard, give me a minute to think it through.

Nicole:

Sarcastic much?

Nicole:

1 attachment
mhchen:

So we are given that "the distance from 5 to 12" is equal to "3x-2" and we are trying to find the value of x. Let A be the sentence "the distance from 5 to 12" Let B be the sentence "3x-2" then we can construct A=B Then "then distance from 5 to 12" = "3x-2" By the definition of distance, the distance between 2 values, a and b, is |a-b|, which is equivalent to |b-a| Therefore the distance from 5 to 12 is |5-12| = |-7| = 7 Thus 7 = 3x - 2 By the additive property of equality, 7+2 = 3x - 2 + 2 Then 9 = 3x By the multiplicative property of equality (1/3)9=(1/3)3x Then 3 = x Hence, our solution, x = 3.

Nicole:

thank you

Nicole:

1 attachment
Nicole:

?

mhchen:

Did you set -3x+1 = 2 - (-2)?

Nicole:

that would be x=-1

mhchen:

ye

Nicole:

1 attachment
mhchen:

3x+2x = 120 right?

Nicole:

Why are we adding?

Nicole:

1 attachment
Nicole:

so this is gonna be 3x+2x=125?

mhchen:

Because ABD + DBC = ABC right? that's why we add

mhchen:

and yes for the 2nd question

Nicole:

okay

1 attachment
Nicole:

?

Nicole:

put them equal to each other?

Nicole:

I mean add?

mhchen:

You know that a straight line is 180 degrees right?

Nicole:

Yes so it will be 6x+120=180?

mhchen:

yes

Nicole:

1 attachment
Nicole:

put them equal to each other?

Nicole:

?

mhchen:

yeah

Nicole:

If lines are perpendicular, what do you know about the angles at which they intersect?

Nicole:

@Vocaloid

dude:

 Perpendicular lines just form a right angle

Nicole:

Correct

Nicole:

what about when they intersect

dude:

There isn't enough information to answer that I'm assuming "If lines are perpendicular, what do you know about the angles at which they intersect?" refers only to perpendicular lines

Nicole:

Yes but im not getting the question to answer it

dude:

I'm confused on what you mean I think the question is only asking about the measure of angles when the intersect perpendicularly (unless this is a whole separate question)

Nicole:

No I just don't know how to answer the question

dude:

Oh You can just define what perpendicular lines are Perpendicular lines are lines that always intersect at a 90º angle

Nicole:

Oh okay got it

Nicole:

If two lines are perpendicular and they intersect at an angle that measures 5x – 12, what is the value of x?

dude:

Well we know that their angle measure is 90º Just set it equal to 90 and solve for x \(5x-12=90\)